simpilify : 1+(3+1)(3^2+1)(3^4+1)(3^16+1)(3^32+1)(3^64+1)=?

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
let x=(3+1)(3^2+1)(3^4+1)(3^16+1)(3^32+1)(3^64+1)
x*(3-1)=(3-1)*(3+1)(3^2+1)(3^4+1)(3^16+1)(3^32+1)(3^64+1)
=(3^2-1)*(3^2+1)(3^4+1)(3^16+1)(3^32+1)(3^64+1)
=(3^4-1)*(3^4+1)(3^16+1)(3^32+1)(3^64+1)
........
=3^128-1
or, x+1= 3^128